OpenStudy (alprincenofl):

Use mathematical induction to show that 1+2^1+2^2+⋯+2^n=2^(n+1)-1

OpenStudy (perl):

is it true for n = 1

OpenStudy (alprincenofl):

http://prntscr.com/57tcdd

OpenStudy (alprincenofl):

@perl it's not a true or false, it's "show that"

OpenStudy (perl):

yes, that is part of the induction proof though

OpenStudy (perl):

'basis case'

OpenStudy (perl):

we want to prove that 2^0+2^1+2^2+...+2^n=2^(n+1)-1 , is true for all n >= 0 basis case; statement true for n=0 ? inductive step: if true for n=k, show it is true for n=k+1

OpenStudy (perl):

ok so far?

OpenStudy (alprincenofl):

Yes clear, I am listening

OpenStudy (perl):

We want to prove that 2^0+2^1+2^2+...+2^n=2^(n+1)-1 , is true for all n >= 0 basis case; statement true for n=0 ? inductive step: if true for n=k, show it is true for n=k+1 We want to prove that 2^0+2^1+2^2+...+2^n=2^(n+1)-1 , is true for all n >= 0 step 1: basis case; statement true for n=0 ? 2^0 = 2^(0+1) -1 1 = 2-1 1=1 This is true (check). Step 2: inductive case: Prove 'if true for n=k, show it is true for n=k+1' So we are given: 2^0 + 2^1 + ... + 2^k = 2^(k+1) - 1 add 2^(k+1) to both sides 2^0 + 2^1 + ... + 2^k + 2^(k+1)= 2^(k+1) - 1 + 2^(k+1) (this is true since equals added to equals are equal) Now simplify the right side of that , using properties of exponents. 2^0 + 2^1 + ... + 2^k + 2^(k+1)= 2^(k+1) + 2^(k+1) - 1 2^0 + 2^1 + ... + 2^k + 2^(k+1)= 2*2^(k+1) - 1 2^0 + 2^1 + ... + 2^k + 2^(k+1)= 2^(k+2) - 1 And this is what we wanted to prove, P(k) -> P(k+1)

OpenStudy (alprincenofl):

@perl much appreciated!